OpenStudy (anonymous):

WILL REWARD!! Simplify the trigonometric expression. sin^2 theta / 1-cos theta A. 1+cos theta B. sin theta C. 1-sin theta / cos theta D. 1+sin theta / cos theta

OpenStudy (luigi0210):

Use the identity: \(\Large sin^2 x+cos^2 x=1\)

OpenStudy (anonymous):

And then what?

OpenStudy (luigi0210):

Well when you get what you need, substitute back in and you'll get a simplified answer.

OpenStudy (anonymous):

So I have sin^2theta+1-cos^2theta=1

OpenStudy (anonymous):

solve this \(\left(\sin^2\theta +\cos^2\theta = 1\right)\) for \(\sin^2\theta\) and sub into your expression